# ExaModelsPower.jl: A GPU-Compatible Modeling Library for Nonlinear Power System Optimization

### v1

@@ IV-D GOC3 Formulation
$10^{-4}$ MadNLP (LK, G) 4 5294.44 7.167e-4
MadNLP (HK, G) 1 10758.4 7.745e-9
MadNCL (G) 0 - -
Ipopt (C) 1 18689.26 1.954e-7
MadNLP (LK, C) 1 15667.51 4.728e-8
To address the limitations of GPU solvers at higher tolerance, we also provide the ability to warm start the GOC3 model using an existing solution. The availability of the full model sheds light on the solution quality of existing results. When running on the unit commitment solutions generated by the benchmark solver, objective values 0.08% greater than the benchmark were achieved. The best known solutions were still $\approx$0.005% better than those found using ExaModelsPower.jl, but generating the GOC3 model using the best known UC solutions would likely yield new best known solutions to the full problem. Objective quality results were generated separately from the benchmarking results shown earlier.
@@ IV-E Break-Down of Total Runtime
This section analyzes the total runtime time (time between initial call and reported solution). The total runtime can be broken down into model construction time, solver initialization time, AD time, linear solve time, and solver internal time. Model construction time is the time spent outside the solver (e.g., building the JuMP model or ExaModel); solver initialization time is the time spent in the solver before the first iteration (e.g., aperforming symbolic factorization); AD time is the time spent computing derivatives; linear solve time is the time spent solving linear systems; and solver internal time is the remainder (e.g., computing step directions). The results in Sections IV-B, IV-C and IV-D reports the time excluding model construction time, which is the common practice in solver benchmarking. When the same problem is solved again with the same sparsity pattern (e.g., no switching), the model construction time and solver initialization time can be eliminated.
@@
PowerModels-CPU(918.5s)ExaModels-CPU(393.5s)ExaModels-GPU(16.1s)00.20.20.40.40.60.60.80.811Solver Configuration (Runtime [s])Runtime ProportionConstructionInitializationInternalADLinear Solve
@@ V Conclusions
This paper introduces ExaModelsPower.jl, a GPU-enabled open-source modeling library for power system optimization. This library provides a model implementation for various power system optimization problems, focusing on large NLP formulations, such as multi-period OPF with and without storage, and the GOC3 security-constrained formulation. Our benchmarking results demonstrate that when solved up to moderate precision, GPU solvers can achieve significant speedups compared to CPU solvers, especially for large-scale problems. This highlights the potential of GPU-accelerated optimization for addressing complex power system problems that were previously computationally prohibitive. However, none of the tested solvers could solve all instances of the GOC3 formulation up to $10^{-8}$ precision. This underscores the need for further research into alternative problem formulations that enable practical solution times for security-constrained multi-period AC OPF at scale.

### v2

@@ IV-D GOC3 Formulation
$10^{-4}$ MadNLP (LK, GPU) 4 5294.44 7.167e-4
MadNLP (HK, GPU) 1 10758.4 7.745e-9
MadNCL (GPU) 0 - -
Ipopt (CPU) 1 18689.26 1.954e-7
MadNLP (LK, CPU) 1 15667.51 4.728e-8
Cvio = maximum constraint violation.
To address the limitations of GPU solvers at higher tolerance, we also provide the ability to warm start the GOC3 model using an existing solution. The availability of the full model sheds light on the solution quality of existing results. When running on the unit commitment solutions generated by the benchmark solver (nvar 140k-330k), objective values 0.08% greater than the benchmark were achieved. The best known solutions were still $\approx$0.005% better than those found using ExaModelsPower.jl, but generating the GOC3 model using the best known UC solutions would likely yield new best known solutions to the full problem. Objective quality results were generated separately from the benchmarking results shown earlier.
@@ IV-E Break-Down of Total Runtime
This section analyzes the total runtime time (time between initial call and reported solution) to solve the 78,484 node epigrids ACOPF case. The total runtime can be broken down into model construction time, solver initialization time, AD time, linear solve time, and solver internal time. Model construction time is the time spent outside the solver (e.g., building the JuMP model or ExaModel); solver initialization time is the time spent in the solver before the first iteration (e.g., performing symbolic factorization); AD time is the time spent computing derivatives; linear solve time is the time spent solving linear systems; and solver internal time is the remainder (e.g., computing step directions). The results in Sections IV-B, IV-C and IV-D reports the time excluding model construction time, which is the common practice in solver benchmarking. When the same problem is solved again with the same sparsity pattern (e.g., no switching), the model construction time and solver initialization time can be eliminated.
@@
PowerModels-CPU(918.5s)ExaModelsPower-CPU(393.5s)ExaModelsPower-GPU(16.1s)00.20.20.40.40.60.60.80.811Solver Configuration (Runtime [s])Runtime ProportionConstructionInitializationInternalADLinear Solve
@@ V Conclusions
This paper introduces ExaModelsPower.jl, a GPU-enabled open-source modeling library for power system optimization. This library provides a model implementation for various power system optimization problems, focusing on large NLP formulations, such as multi-period OPF with and without storage, and the GOC3 security-constrained formulation. Our benchmarking results demonstrate that when solved up to moderate precision, GPU solvers can achieve significant speedups compared to CPU solvers, especially for large-scale problems. This highlights the potential of GPU-accelerated optimization for addressing complex power system problems that were previously computationally prohibitive. However, none of the tested solvers could solve all instances of the GOC3 formulation up to $10^{-8}$ precision. This underscores the need for further research into alternative problem formulations that enable practical solution times for security-constrained multi-period AC OPF at scale and understanding the tradeoffs between such formulations. Future work can also augment ExaModelsPower.jl to solve numerous ACOPF problems in batch.